OpenStudy (anonymous):

how do you solve this x/4+7=10?? help!

OpenStudy (mertsj):

Subtract 7 from both sides and then multiply by 4 on both sides.

OpenStudy (saifoo.khan):

\[\frac x4 + 7 = 10\]Subtract "7" from both sides, \[\frac x4 = 10 - 7\] \[\frac x4 = 3\] Multiply 4 on both sides, \[x = 4 \times 3\] \[x = 12\]

OpenStudy (anonymous):

thx guyz!!

OpenStudy (anonymous):

how do you solve x+28=40?

OpenStudy (anonymous):

revers oporation